# CU-JADE: A Method for Traversing Extinction Jumps along the Line-of-Sight

### v1

@@ 5.2.2 Cygnus Region
In previous work (Zhang et al., 2024), we have identified MCs in the foreground layers located on the Cygnus Rift. In Figure 8, we cross-verify the extinction structures and those $\rm{}^{13}CO$ traced MCs in the foreground ($\lesssim 1.1$ kpc). Using the multi-jump detection method, we again confirm significant extinction in the Cygnus Rift, which is also distributed over various distance layers, for example, structures at $\sim 1$ kpc (deep red), the 800 pc loop (light red and dashed ellipse), and closer MCs likely associated with Cyg OB7 (structures with even lighter colors below the 800 pc loop at $82^{\circ}\lesssim l\lesssim 87^{\circ}$, $-4^{\circ}\lesssim b\lesssim-2^{\circ}$). We can see that strong extinction structures ($\gtrsim 1$ mag, see black contours in Figure 8) match very well with the molecular gas traced by $\rm{}^{13}CO$ (see white contours in Figure 8).
@@
Figure 10: For the case close to the projected edge of the Cygnus Rift, four examples showing an anti-correlation between extinction and integrated CO intensity in dense regions, which can also serve as strong evidence for determining distances. Here, the greyscale image represents the integrated intensity of MWISP $\rm{}^{12}CO$. Similarly, thin red contours indicate the cumulative values of the extinction within the corresponding distance interval. Blue contours denote local minima of extinction (identified using $minimum\_filter$ in $Scipy.ndimage$) that coincide with dense gas regions ($\rm\gtrsim 20~{}K~{}km~{}s^{-1}$). The distance interval, velocity interval, and determined distances are annotated at the top of the figure.
@@
In more complex regions, the integrated intensity of the CO emission is also positively correlated with the extinction within the certain distance interval for the whole dense gas environments (see Figure 10). Intriguingly, in the brightest CO emission region of the four MCs, the integrated extinction in the distance measuring interval is smaller or empty. It indicates an anti-correlation between CO brightness and the extinction in the highest column density region. The dense part in MCs itself creates an “extinction wall” (Straizys et al., 1993; Zhang et al., 2024) that obscures background stars, preventing the detection of the extinction components. This effect could also be revealed in the $D$-$A$ map. In Figure 10, these regions in column density-extinction anti-correlation are highlighted with blue contours, agreeing with dense CO emission at local minima in extinction.
@@
In the CygX-South region, some MCs with enhanced CO emission are associated with bright mid-infrared features (e.g., cloud 5 in Figure 10, or Clouds A, B in Schneider et al., 2006, 2007). We also identify more mid-infrared bright MCs in Zhang et al. (2024) and the other three MCs (cloud 6, 7, and 8 in Figure 10) in this work. These MCs near photodissociation regions are suggested to be affected by the stellar winds and radiation from Cyg OB1 and OB2 (Schneider et al., 2006, 2007), featuring a head-to-tail structure pointing toward the stars association with bright CO emission. Through distance measurements with accurate Gaia parallaxes, for the first time, we confirm that these cometary MCs are located at a distance of $\sim 1.7$ kpc within an uncertainty of 10% – 15%. These results strongly suggest that these interesting clouds are physically associated with the nearby HMSFRs (i.e., main clusters of Cyg OB1 and OB2 at $\sim 1.7-1.8$ kpc; see Berlanas et al., 2019; Orellana et al., 2021; Quintana & Wright, 2021).
@@
Unfortunately, in regions with particularly high column densities (e.g., see details in Figures 8, 9, and 10 in Zhang et al., 2024), star sampling in $D$-$A$ map is incomplete due to intense extinction of multiple gas layers. It is difficult to measure the dense gas structure behind the Cygnus Rift based only on the Gaia parallax data and visible V/G band extinction. Such scenarios could also be seen toward Aquila Rift region (see Figure 5), where the number of stars drop suddenly at a certain distance. In the future, we will combine data from the near-infrared surveys (e.g., Kollmeier et al., 2017) and new models (e.g., Rezaei Kh. et al., 2018) to explore the molecular gas distribution behind the “extinction wall”.
@@
We present the first application of the CUSUM-based jump point analysis method, named CU-JADE, to identify multiple extinction jumps along the LoS. Our results demonstrate CU-JADE’s capability to detect weak extinction jump points (where $\Delta A_{\lambda}\approx\sigma$) in $D$-$A$ diagrams, even in complex Galactic disk regions with overlapping gas features. The method iteratively analyzes $D$-$A$ diagrams using CUSUM statistics, focusing on cumulative extinction increments rather than traditional curve fitting for $D$-$A$ data series. Through mathematical derivation and numerical simulations, we established the relationships between $\Delta A_{\lambda}$, $\sigma$, and $CL$, while evaluating the impacts of sample sparsity (low stellar density $n$) and diffuse gas contributions (i.e., the linear slope $k$ in $D$-$A$ diagrams).
@@
Applied to mock data, CU-JADE shows excellent accuracy and completeness in distance measurements. Validation against HMSFR maser parallaxes revealed no systematic offsets (see Figure 3). Using SHEDR3 data, we constructed an all-sky 3D extinction map within 4 kpc (Figures 4 and 5), whose dense components correlate well with YOCs and masers. The map reveals known structures (e.g., Split, Radcliffe Wave, Local Bubble clouds) while providing new insights into their 3D distribution. Comparison of our findings with the ZGR23 dataset reveals good agreement in local structures, while also identifying previously unresolved features in regions toward the outer Galaxy and the Galactic center. Notably, our results exhibit a reduced “Finger of God” effect in top-down projections. By optimizing parameter adjustments and selecting stellar samples, CU-JADE’s effectiveness and capabilities can be further enhanced.
@@
We conducted detailed case studies on two typical regions, Cepheus and Cygnus by using the CU-JADE. For Cepheus in relatively high Galactic latitude, we have reconstructed a multilayer extinction structures at different distances: the nearby MCs with $\sim 0$ $\rm km~{}s^{-1}$ at $\sim 300$ pc distributed throughout the whole region, intermediate-distance clouds with $\sim-12$ $\rm km~{}s^{-1}$ at $\sim 700-1000$ pc extend from the Galactic plane to $b\sim 15^{\circ}$, and Perseus Arm clouds at $\sim$ $2-3$ kpc (Figure 7).
@@
In the more complicated Cygnus region near the Galactic plane, CU-JADE successfully delineates both the Cygnus Rift MCs within 1 kpc (Figure 8) and more distant MC structures at $1-3$ kpc (Figures 9 and 10), enabling the first distance measurements to MCs associated with Cyg OB2/OB1 through correlation with MWISP CO data. Being able to measure distances to MCs associated with massive stars is a big deal because OB stars are short-lived, indicating regions of recent or ongoing star formation. This correlation shows that CU-JADE can handle more challenging environments.
@@
The CU-JADE method shows exceptional adaptability across spatial scales (a few arcmins to several tens of arcmins) and extinction environments, enabling robust MC studies at various distances. Our 3D extinction mapping technique premits parameter optimization to align with both in CO-traced MCs and CO-dark MCs, accounting for stellar sampling density and local extinction conditions. Validation confirms the method’s effectiveness for both nearby ($\lesssim 1$ kpc) and more distant ($1-3$ kpc) structures. Future work will focus on three main areas: (1) determining precise distances for MWISP CO-traced MCs using multi-band data, (2) investigating structures related to CO-dark and/or CO-faint MCs, and (3) refining parameter space exploration to enhance the fidelity of maps in complex Galactic regions. This approach offers a powerful tool for linking dust extinction features with the morphological characteristics of MCs in the ISM.
@@
In our current work, we find that a single parameter value “$C$” is sufficient for generating reliable thresholds across the entire sky to identify significant extinction jump structures. This simplified approach is justified given our limited prior knowledge of dust distribution, while minimizing the model’s dependence on the output. Consequently, Mode I (Section 3.2.2) represents a more conservative strategy for detecting extinction jumps, as it disregards variation in both source density and extinction properties along different sightlines (e.g., spatial fluctuations in extinction curve parameters and stellar densities).
@@
Future studies will focus on three key improvements to the dust map: (1) Optimization of Mode II through implementation of adaptive scaling factors to correct for spatial inhomogeneities in stellar density distributions. (2) Minimization of inherent projection effects in our sampling grid using HEALPix-based resolution methods (Górski et al., 2005). (3) Investigation of alternative grid classifications to recover local extinction structures (e.g., $\lesssim 500$ pc), employing a progressive analysis strategy from high-Galactic-latitude regions ($|b|\gtrsim 30^{\circ}$) to the Galactic disk. Additionally, independent gas tracers (e.g., HI and CO) will provide critical validation for both identifying and compensating for potential systematic biases throughout the optimization procedure.
@@
Acknowledgment. We gratefully acknowledge the referee’s expert feedback, which significantly enhanced the paper’s methodological robustness and interpretive depth. This research made use of the data from the Milky Way Imaging Scroll Painting (MWISP) project, which is a multiline survey in $\rm{}^{12}CO$/$\rm{}^{13}CO$/$\rm C^{18}O$ along the northern Galactic plane with the PMO 13.7 m telescope. We are grateful to all the members of the MWISP working group, particularly the staff members at the PMO-13.7m telescope, for their long-term support. MWISP was sponsored by the National Key R&D Program of China with grants 2023YFA1608000, 2017YFA0402701, and the CAS Key Research Program of Frontier Sciences with grant QYZDJ-SSW-SLH047. This work is supported by NSFC grants 12173090. X.C. acknowledges the support from the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the support by the CAS International Cooperation Program (grant No. 114332KYSB20190009). This work also made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The online catalogs of all-sky LoS extinction jumps, along with the CUSUM-based algorithm scripts supporting the findings of this study, are available at ScienceDB: 10.57760/sciencedb.24650.
Appendix A Complications of Multi-Jump Points Detection
@@ A.1 The CUSUM statistics
$D$ is the distance set in an ascending order, $A_{i}^{j}$ is the $i$-th $A_{\lambda}$ value of $j$-th group. $\mu_{j}$ and $\sigma_{j}$ are the mean and dispersion of $A_{\lambda}$ in $j$-th group. $d_{min}$ and $d_{max}$ denote the minimum and maximum distances of the samples, respectively. Additionally, $d_{j}$ is the distance corresponding to the $j$-th jump point. In our work, the mean values $\mu_{1}$, $\mu_{2}$, $\cdots$, $\mu_{j}$, $\mu_{j+1}$ are ordered such that $\mu_{1}<\mu_{2}<\cdots<\mu_{j}<\mu_{j+1}$. And similarly, for most cases, $\sigma_{1}\leq\sigma_{2}\leq\cdots\leq\sigma_{j}\leq\sigma_{j+1}$. As with Equation (1), the mean $A_{\lambda}$ value of all samples is
@@
$N$ represents the number of samples. The $D$-$S$ plot exhibits $j$ turning points, and the spike is often located at the midpoints of these turning points, which is initially identified as a change point (see Figure 1). It should be noted that the spike does not necessarily indicate the most prominent change point. However, the mean $A_{\lambda}$ values on both sides are closest to $\bar{A}$.
The calculation of $S^{0}_{\rm diff}$ for the $D$-$S$ curve with multiple jump points is not easily derived directly from the mathematical expression. We thus simplify the problem as follows: During iterations, it happens that a subset contains only two sets of $A_{\lambda}$ values with one extinction jump. The mean values of the two groups are $\mu_{j}$ and $\mu_{j+1}$, respectively, with the number of observations in each group being $N_{j}$ and $N_{j+1}$. To simplify the analysis of cumulative data, fluctuations caused by the dispersion of $A_{\lambda}$ are canceled out. The prominence of the CUSUM statistic $S$ is
@@
In addition to those factors discussed in Section 3.1, the other scenarios would also affect the multiple extinction-jump detection. For example, the closer the jump point is to the middle of the sequence, the larger the $S^{0}_{\rm diff}$ value will be. The same situation also appears when the more subsequent jump points are detected (e.g., from 1 to 3 in Table 1). In other words, the first jump point at the middle sequence is more likely to be a false positive. This scenario might affect all of our traversed jumps and can be identified using numerical methods. However, the overall impact of these factors is minor, which can be further confirmed by simulations (see Section 4.1 and Table 1).
Figure A1: An example for simple one-jump-point case, denoting the relationship between $\rm S_{diff}^{Avg}$, $\Delta A_{G}$, $\sigma$, and $n$. (a) Set stars number $N$=1000, $S^{\rm avg}_{\rm diff}\propto\sigma$ when $\sigma\gtrsim\Delta A$. (b) $N$=1000, $S^{\rm avg}_{\rm diff}\propto\Delta A$ when $\Delta A\gg\sigma$. (c) $S^{\rm avg}_{\rm diff}\propto\sqrt{n}$. The zoom-in area is presented for comparable $\Delta A_{G}$ and $\sigma$.

### v2

@@ 5.2.2 Cygnus Region
In our previous work (Zhang et al., 2024), we have identified MCs in the foreground layers located on the Cygnus Rift. In Figure 8, we cross-verify the extinction structures and those $\rm{}^{13}CO$ traced MCs in the foreground ($\lesssim 1.1$ kpc). Using the multijump detection method, we again confirm significant extinction in the Cygnus Rift, which is also distributed over various distance layers, e.g., structures at $\sim 1$ kpc (deep red), the 800 pc loop (light red and dashed ellipse), and closer MCs likely associated with Cyg OB7 (structures with even lighter colors below the 800 pc loop at $82^{\circ}\lesssim l\lesssim 87^{\circ}$, $4^{\circ}\lesssim b\lesssim-2^{\circ}$). We can see that strong extinction structures ($\gtrsim 1$ mag, see black contours in Figure 8) match very well with the molecular gas traced by $\rm{}^{13}CO$ (see white contours in Figure 8).
@@
Figure 10: For the case close to the projected edge of the Cygnus Rift, four examples showing an anticorrelation between extinction and integrated CO intensity in dense regions, which can also serve as strong evidence for determining distances. Here, the greyscale image represents the integrated intensity of MWISP $\rm{}^{12}CO$. Similarly, thin red contours indicate the cumulative values of the extinction within the corresponding distance interval. Blue contours denote local minima of extinction (identified using minimum_filterminimum\_filter in $Scipy.ndimage$) that coincide with dense gas regions ($\rm\gtrsim 20~K~km~s^{-1}$). The distance interval, velocity interval, and determined distances are annotated at the top of the figure.
@@
In more complex regions, the integrated intensity of the CO emission is also positively correlated with the extinction within the certain distance interval for the whole dense gas environments (see Figure 10). Intriguingly, in the brightest CO emission region of the four MCs, the integrated extinction in the distance measuring interval is smaller or empty. It indicates an anticorrelation between CO brightness and the extinction in the highest column density region. The dense part in MCs itself creates an “extinction wall” (Straizys et al., 1993; Zhang et al., 2024) that obscures background stars, preventing the detection of the extinction components. This effect could also be revealed in the $D$-AA map. In Figure 10, these regions in column density-extinction anticorrelation are highlighted with blue contours, agreeing with dense CO emission at local minima in extinction.
@@
In the CygX-South region, some MCs with enhanced CO emission are associated with bright mid-infrared features (e.g., cloud 5 in Figure 10, or Clouds A and B in Schneider et al., 2006, 2007). We also identify more mid-infrared bright MCs in Zhang et al. (2024) and the other three MCs (clouds 6, 7, and 8 in Figure 10) in this work. These MCs near photodissociation regions are suggested to be affected by the stellar winds and radiation from Cyg OB1 and OB2 (Schneider et al., 2006, 2007), featuring a head-to-tail structure pointing toward the stars association with bright CO emission. Through distance measurements with accurate Gaia parallaxes, for the first time, we confirm that these cometary MCs are located at a distance of $\sim 1.7$ kpc within an uncertainty of 10% – 15%. These results strongly suggest that these interesting clouds are physically associated with the nearby HMSFRs (i.e., main clusters of Cyg OB1 and OB2 at $\sim 1.7-1.8$ kpc; see Berlanas et al., 2019; Orellana et al., 2021; Quintana & Wright, 2021).
@@
Unfortunately, in regions with particularly high column densities (e.g., see details in Figures 8, 9, and 10 in Zhang et al., 2024), star sampling in the $D$-AA map is incomplete due to intense extinction of multiple gas layers. It is difficult to measure the dense gas structure behind the Cygnus Rift based only on the Gaia parallax data and visible VV/$G$-band extinction. Such scenarios could also be seen toward Aquila Rift region (see Figure 5), where the number of stars drop suddenly at a certain distance. In the future, we will combine data from the near-infrared surveys (e.g., Kollmeier et al., 2017) and new models (e.g., Rezaei Kh. et al., 2018) to explore the molecular gas distribution behind the “extinction wall”.
@@
We present the first application of the CUSUM-based jump point analysis method, named CU-JADE, to identify multiple extinction jumps along the LoS. Our results demonstrate CU-JADE’s capability to detect weak extinction-jump points (where $\Delta A_{\lambda}\approx\sigma$) in $D$-AA diagrams, even in complex Galactic disk regions with overlapping gas features. The method iteratively analyzes $D$-AA diagrams using CUSUM statistics, focusing on cumulative extinction increments rather than traditional curve fitting for $D$-AA data series. Through mathematical derivation and numerical simulations, we established the relationships between $\Delta A_{\lambda}$, $\sigma$, and $CL$, while evaluating the impacts of sample sparsity (low stellar density $n$) and diffuse gas contributions (i.e., the linear slope $k$ in $D$-AA diagrams).
@@
Applied to mock data, CU-JADE shows excellent accuracy and completeness in distance measurements. Validation against HMSFR maser parallaxes revealed no systematic offsets (see Figure 3). Using SHEDR3 data, we constructed an all-sky 3D extinction map within 4 kpc (Figures 4 and 5), whose dense components correlate well with YOCs and masers. The map reveals known structures (e.g., Split, Radcliffe Wave, and Local Bubble clouds) while providing new insights into their 3D distribution. Comparison of our findings with the ZGR23 dataset reveals good agreement in local structures, while also identifying previously unresolved features in regions toward the outer Galaxy and the Galactic center. Notably, our results exhibit a reduced “Finger of God” effect in top-down projections. By optimizing parameter adjustments and selecting stellar samples, CU-JADE’s effectiveness and capabilities can be further enhanced.
@@
We conducted detailed case studies on two typical regions, Cepheus and Cygnus, by using the CU-JADE. For Cepheus in relatively high Galactic latitude, we have reconstructed a multilayer extinction structures at different distances: the nearby MCs with $\sim 0$ $\rm km~s^{-1}$ at $\sim 300$ pc distributed throughout the whole region, intermediate-distance clouds with $\sim-12$ $\rm km~s^{-1}$ at $\sim 700-1000$ pc extend from the Galactic plane to $b\sim 15^{\circ}$, and Perseus Arm clouds at $\sim$ 232-3 kpc (Figure 7).
@@
In the more complicated Cygnus region near the Galactic plane, CU-JADE successfully delineates both the Cygnus Rift MCs within 1 kpc (Figure 8) and more distant MC structures at 131-3 kpc (Figures 9 and 10), enabling the first distance measurements to MCs associated with Cyg OB2/OB1 through correlation with MWISP CO data. Being able to measure distances to MCs associated with massive stars is a big deal because OB stars are short-lived, indicating regions of recent or ongoing star formation. This correlation shows that CU-JADE can handle more challenging environments.
@@
The CU-JADE method shows exceptional adaptability across spatial scales (a few arcmins to several tens of arcmins) and extinction environments, enabling robust MC studies at various distances. Our 3D extinction mapping technique permits parameter optimization to align with both in CO-traced MCs and CO-dark MCs, accounting for stellar sampling density and local extinction conditions. Validation confirms the method’s effectiveness for both nearby ($\lesssim 1$ kpc) and more distant (131-3 kpc) structures. Future work will focus on three main areas: (1) determining precise distances for MWISP CO-traced MCs using multiband data, (2) investigating structures related to CO-dark and/or CO-faint MCs, and (3) refining parameter space exploration to enhance the fidelity of maps in complex Galactic regions. This approach offers a powerful tool for linking dust extinction features with the morphological characteristics of MCs in the ISM.
@@
In our current work, we find that a single parameter value “CC” is sufficient for generating reliable thresholds across the entire sky to identify significant extinction-jump structures. This simplified approach is justified given our limited prior knowledge of dust distribution, while minimizing the model’s dependence on the output. Consequently, Mode I (Section 3.2.2) represents a more conservative strategy for detecting extinction jumps, since it disregards variation in both source density and extinction properties along different sightlines (e.g., spatial fluctuations in extinction curve parameters and stellar densities).
@@
Future studies will focus on three key improvements to the dust map: (1) optimization of Mode II through implementation of adaptive scaling factors to correct for spatial inhomogeneities in stellar-density distributions, (2) minimization of inherent projection effects in our sampling grid using HEALPix-based resolution methods (Górski et al., 2005), and (3) Investigation of alternative grid classifications to recover local extinction structures (e.g., $\lesssim 500$ pc), employing a progressive analysis strategy from high Galactic latitude regions ($|b|\gtrsim 30^{\circ}$) to the Galactic disk. Additionally, independent gas tracers (e.g., HI and CO) will provide critical validation for both identifying and compensating for potential systematic biases throughout the optimization procedure.
@@
Acknowledgment. We gratefully acknowledge the referee’s expert feedback, which significantly enhanced the paper’s methodological robustness and interpretive depth. This research made use of the data from the Milky Way Imaging Scroll Painting (MWISP) project, which is a multiline survey in $\rm{}^{12}CO$$\rm{}^{13}CO$$\rm C^{18}O$ along the northern Galactic plane with the PMO 13.7 m telescope. We are grateful to all the members of the MWISP working group, particularly the staff members at the PMO-13.7m telescope, for their long-term support. MWISP was sponsored by the National Key R&D Program of China with grants 2023YFA1608000, 2017YFA0402701, and the CAS Key Research Program of Frontier Sciences with grant QYZDJ-SSW-SLH047. This work is supported by NSFC grants 12173090. X.C. acknowledges the support from the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the support by the CAS International Cooperation Program (grant No. 114332KYSB20190009). This work also made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
The online catalogs of all-sky LoS extinction jumps, along with the CUSUM-based algorithm scripts supporting the findings of this study, are available at ScienceDB: 10.57760/sciencedb.24650.
Appendix A Complications of Multijump Points Detection
@@ A.1 The CUSUM statistics
$D$ is the distance set in an ascending order, $A_{i}^{j}$ is the $i$th $A_{\lambda}$ value of $j$th group. $\mu_{j}$ and $\sigma_{j}$ are the mean and dispersion of $A_{\lambda}$ in $j$th group. $d_{min}$ and $d_{max}$ denote the minimum and maximum distances of the samples, respectively. Additionally, $d_{j}$ is the distance corresponding to the $j$th jump point. In our work, the mean values $\mu_{1}$, $\mu_{2}$, $\cdots$, $\mu_{j}$, $\mu_{j+1}$ are ordered such that $\mu_{1}<\mu_{2}<\cdots<\mu_{j}<\mu_{j+1}$. Similarly, for most cases, $\sigma_{1}\leq\sigma_{2}\leq\cdots\leq\sigma_{j}\leq\sigma_{j+1}$. As with Equation (1), the mean $A_{\lambda}$ value of all samples is
@@
NN represents the number of samples. The $D$-$S$ plot exhibits $j$ turning points, and the spike is often located at the midpoints of these turning points, which is initially identified as a change point (see Figure 1). It should be noted that the spike does not necessarily indicate the most prominent change point. However, the mean $A_{\lambda}$ values on both sides are closest to $\bar{A}$.
The calculation of $S^{0}_{\rm diff}$ for the $D$-$S$ curve with multiple jump points is not easily derived directly from the mathematical expression. We thus simplify the problem as follows. During iterations, it happens that a subset contains only two sets of $A_{\lambda}$ values with one extinction jump. The mean values of the two groups are $\mu_{j}$ and $\mu_{j+1}$, respectively, with the number of observations in each group being $N_{j}$ and $N_{j+1}$. To simplify the analysis of cumulative data, fluctuations caused by the dispersion of $A_{\lambda}$ are canceled out. The prominence of the CUSUM statistic $S$ is
@@
In addition to those factors discussed in Section 3.1, the other scenarios would also affect the multiple extinction-jump detection. For example, the closer the jump point is to the middle of the sequence, the larger the $S^{0}_{\rm diff}$ value will be. The same situation also appears when the more subsequent jump points are detected (e.g., from 1 to 3 in Table 1). In other words, the first jump point at the middle sequence is more likely to be an $FP$. This scenario might affect all of our traversed jumps and can be identified using numerical methods. However, the overall impact of these factors is minor, which can be further confirmed by simulations (see Section 4.1 and Table 1).
Figure A1: An example for simple one-jump-point case, denoting the relationship between $\rm S_{diff}^{Avg}$, $\Delta A_{G}$, $\sigma$, and $n$. (a) Set stars number NN=1000, $S^{\rm avg}_{\rm diff}\propto\sigma$ when $\sigma\gtrsim\Delta A$. (b) NN=1000, $S^{\rm avg}_{\rm diff}\propto\Delta A$ when $\Delta A\gg\sigma$. (c) $S^{\rm avg}_{\rm diff}\propto\sqrt{n}$. The zoom-in area is presented for comparable $\Delta A_{G}$ and $\sigma$.